OpenStudy (anonymous):

Suppose that the dollar value v(t) of a certain car that is t years old is given by the following exponential function. v(t) = 27,500 (0.76)^t -find the initial value of the car - does the funtion represent growth or decay? - by what percent does the value of the car change each year?

OpenStudy (anonymous):

@mathsciencehistory

OpenStudy (anonymous):

umm idk actually

OpenStudy (anonymous):

@Gummibear127

OpenStudy (anonymous):

I just realized when I posted the question the form of it got all mixed up idk if it helps at all but I just fixed it haha

OpenStudy (amistre64):

what is the value of time when you start to measure things?

OpenStudy (anonymous):

i just ran it through mathway but all it gave me was 27.is that correct?

OpenStudy (anonymous):

a year?

OpenStudy (amistre64):

if i want to see how long i can hold my breath, i dont start the timer at 1 minute :) t = 0 for an initial value

OpenStudy (amistre64):

now to see if we are increasing or decreasing, what is the change in value between 0t=0 and t=1?

OpenStudy (anonymous):

ok so really quick t=0 so that makes my first answer 20,900 right? or are we still working on that one

OpenStudy (amistre64):

27,500 (0.76)^t .76^0 = 1 soo, 27500*1 = 27500

OpenStudy (amistre64):

for the second one we get an increase or decrease in value when t=1? .76^1 = .76 soo 27500 * .76 = 20900 20900 is less then 27500, so this is a _________ in value over time.

OpenStudy (anonymous):

oh ok decrease

OpenStudy (anonymous):

decay

OpenStudy (amistre64):

now we just need to determine the percent of decrease ... the rate of decay. any thoughts on how to do that one?

OpenStudy (anonymous):

ehh I'm sure it's easy but I didn't realize how easy the first half was until you explained it either lol

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

does it have something to do with the 0.76 in parenthesis

OpenStudy (amistre64):

well, if the site doesnt go down ... it does have something to do with that, but theres a more fundamental approach we can use. we want to compare the change in amount, to the original amount (27500(.76) - 27500(1)) / 27500 we can factor out the 27500, and this reduces to (.76-1) 27500/27500 or simply - (1-.76) ^^ so it does have something with the .76 and the negative in front tells us its decreaseing at a rate of 1-.76

OpenStudy (amistre64):

if we want to get fancy, we start observing properties of this setup ... P(k)^t the original amount is always in front: P the rate, r, can be determined as 1 + r = k, or simply r = k-1 if r is negative, we are decaying, if its positive, we are increasing (caying?)

OpenStudy (anonymous):

haha =) so 76%?